Assuming an efficiency of 38.30%,

38.30%, calculate the actual yield of magnesium nitrate formed from 114.4 g

114.4 g of magnesium and excess copper(II) nitrate.


Mg+Cu(NO3)2⟶Mg(NO)2+Cu



1
Expert's answer
2021-03-09T06:06:25-0500

Mg + Cu(NO3)2 = Mg(NO3)2 + Cu

Mg = Mg(NO3)2

24 g Mg = 148g Mg(NO3)2

1g Mg = 148/24 Mg(NO3)2

114.4g Mg = 148× 114.4g/ 24 Mg(NO3)2

= 705.5g ( This is the theoretical yield)


%efficiency = actual yield/ theorectical yield × 100%

actual yield = % efficiency × theoretical yield/ 100%

= 38.8% ×705.5 /100%

= 270.2g